I'm having problems with writing an equation of a line that is perpendicular because our text book doesn't do a good job explaining it.
"write the equation of a line that is perpendicular to the given line and that passes through the given point." -4x-8y=24; (4,-1) the opions they give you are
1.) y=2x-9
2.) y=1/2x+6
3.) y=1/2x-9
4.) y=-2x-9
I found the negative reciprocal of 4 which is -1/4 and that's as far as I have gotten. After 13 hours of doing this homework I honestly can't understand what I'm reading.

Answer by Alan3354(69443)

"write the equation of a line that is perpendicular to the given line and that passes through the given point." -4x-8y=24; (4,-1) the opions they give you are
-----------------
1st find the slope of the line. Put the eqn into slope-intercept form (that means solve for y)
-4x-8y=24
-x-2y = 6
2y = -x - 6
y = (-1/2)x - 3
That's the form, y = mx + b. m is the slope, -1/2
----------------
The slope of all lines perpendicular is the neg recip, = +2
-----------------
Then use y = mx + b and the point to find b:
-1 = 2*4 + b
b = -9
That makes the eqn y = 2x - 9